OpenStudy (anonymous):

Which would be the domain in this case? f(x)=[x+2]/[x^2+7x+10] I know that we can factorize the denominator like this: (x+2)(x+5). So the roots of the denominator are -2 and -5. But I can cancel out x+2 from the numerator.... So the domain is all real - {-2,-5} OR all real -{-5} ??

OpenStudy (anonymous):

The numbers that make the denominator 0 are not included in the domain, so its all real numbers except -2 and -5

OpenStudy (anonymous):

I made a oopsy

OpenStudy (anonymous):

This is the same as 1/x+5 so domain is all except -5

OpenStudy (anonymous):

But we can rewrite f(x) this way: \[f(x)=\frac{x+2}{x^2+7x+10}=\frac{x+2}{(x+2)(x+5)}=\frac{1}{x+5}\] Isn't this called a "hole"? Could you give me a resource to check this?

OpenStudy (anonymous):

Yeah thats what i was saying, not sure its called anything, but if you graph it -2 is in the domain

OpenStudy (anonymous):

mm.. What about this? http://www.purplemath.com/modules/grphrtnl4.htm

OpenStudy (anonymous):

Thats kinda interesting..

OpenStudy (anonymous):

But there it says that we have to consider the roots of the denominator of the original function. Which is true?